# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ E.4 New Guarantees for DAgger Variant with Trajectory-wise Annotation
By Lemma 28 and Lemma 29, for each $\pi^{n}$, whether it is Markovian or a first-step mixing of Markovian policies, the following holds:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
@@
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step mixing of $\{\pi^{n}\}_{n=1}^{N}$ with uniform weights. Finally, we conclude the proof by applying Lemma 37. ∎
Theorem 34 (Theorem 27 Restated).
If Algorithm 3 is run with a deterministic expert policy $\pi^{E}$, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set to exponential weight algorithm (see Proposition 38). Then, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
Proof.
The proof closely follows Proposition C.2 in [foster2024behavior], tailored for another DAgger variant. However, in this case, we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
@@
$\ell^{n}(\pi)=\log(1/\pi_{u}(a^{n,*}_{1:H}\parallel s^{n}_{1:H}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right).$
It can be observed that $\ell^{n}$’s are 1-exp-concave. Therefore, implementing $\mathbb{A}$ using the exponential weights algorithm (Proposition 38) ensures that the following bound holds almost surely:
@@
Then, Lemma 39 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(\mathbf{E^{\prime}}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
Finally, the second part of the theorem follows by applying Lemma 33. ∎
@@ Appendix F Auxiliary Results
$\frac{1}{2}\parallel p-q\parallel_{1}\leq D_{H}^{2}(p\parallel q)\leq\parallel p-q\parallel_{1}$
Lemma 36 (Performance Difference Lemma [Kakade2002ApproximatelyOA][ross2010efficient]).
@@
Lemma 37 (Lemma D.2. of  [foster2024behavior]).
@@
Proposition 38 (Proposition 3.1 of [cesa2006prediction]).
Suppose $\{\ell^{n}(u)\}_{n=1}^{N}$ is a sequence of $\eta$-exp-concave functions from $\Delta(\mathcal{X})$ to $\mathbb{R}$. For all $x\in\mathcal{X}$, define the weights $w_{x}^{n-1}$ and probabilities $u^{n}(x)$ as follows:
$w_{x}^{n-1}=e^{-\eta\sum_{i=1}^{n-1}\ell_{i}(e_{x})},\quad u^{n}(x)=\frac{w_{x}^{n-1}}{\sum_{x^{\prime}\in\mathcal{X}}w_{x^{\prime}}^{n-1}},$
where $e_{x}$ is the $x$-th standard basis vector in $\mathbb{R}^{|\mathcal{X}|}$. Then, choosing $u^{n}=\{u^{n}(x)\}_{x\in\mathcal{X}}$ (exponential weights used with learning rate $\eta$) satisfies:
@@
Lemma 39 (Restatement of Lemma A.14 in [foster2021statistical]).
Under the realizbility assumption, where there exists $g_{\star}:=g_{i_{\star}}\in\mathcal{G}$ such that for all $n\in[N]$,
$y^{n}\sim g_{\star}^{n}(\cdot\mid x^{n})\mid x^{n},\mathcal{H}^{n-1},$
where $\mathcal{H}^{n-1}$ denotes all histories at the beginning of round $n$.
Then, for any estimation algorithm and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{n=1}^{N}\mathbb{E}_{n-1}\left[D_{\mathcal{H}}^{2}\left(\hat{g}^{n}(x^{n}),g_{\star}^{n}(x^{n})\right)\right]\leq\sum_{n=1}^{N}\left(\ell_{\log}^{n}(\hat{g}^{n})-\ell_{\log}^{n}(g^{n}_{\star})\right)+2\log(\delta^{-1}).$
where $\ell_{\log}^{n}\left(g\right):=\log\left(1/g\left(y^{n}\mid x^{n}\right)\right)$, and $\mathbb{E}_{n}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{H}^{n}]$.
We have the following well-known lemma for causally-conditioned probability (e.g. [ziebart2010modeling]).
Definition 40.
The causally-conditioned probability of state sequence $s_{1:H}$ given action sequence $a_{1:H-1}$, is defined as
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})$
Lemma 41.
For any Markovian policy $\pi$,
and for any first-step mixing of Markovian policy $\pi_{u}$,
Proof.
Eq. (21) follows by noticing that both sides are equal to
$P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
Eq. (22) follows by noticing that both sides are equal to
$\sum_{\nu}u(\nu)P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$
@@ Appendix G Experiment Details
We compare Warm-Stagger with Behavior Cloning (BC) and Stagger on continuous-control tasks from OpenAI Gym MuJoCo [todorov2012mujoco, brockman2016openai] with episode length $H=1000$.
@@
All experiments were conducted on a Linux workstation equipped with an Intel Core i9 CPU (3.3GHz) and four NVIDIA GeForce RTX 2080 Ti GPUs. Our implementation builds on the publicly available DRIL framework [brantley2019disagreement] (https://github.com/xkianteb/dril), with modifications to support online learning. The continuous control environments used in our experiments are: “HalfCheetahBulletEnv-v0”, “AntBulletEnv-v0”, “Walker2DBulletEnv-v0”, and “HopperBulletEnv-v0”. We include an anonymous link to our implementation here: https://github.com/liyichen1998/Interactive-and-Hybrid-Imitation-Learning-Provably-Beating-Behavior-Cloning.
@@
We use four MuJoCo environments: Ant, Hopper, HalfCheetah, and Walker2D. The expert policy is a deterministic MLP pretrained via TRPO [schulman2015trust, schulman2017proximal], with two hidden layers of size 64.
Learner Architecture.
The learner uses the same MLP architecture as the expert. Following [foster2024behavior], we use a diagonal Gaussian policy:
@@
To evaluate the performance of BC against the number of states annotated, we reveal expert state-action pairs sequentially along expert trajectories until the annotation budget is reached. For Stagger, each round it rolls out the latest policy, samples a state uniformly from the trajectory, queries it for the expert action, and updates immediately.
For Warm-Stagger, we begin with BC and switch to Stagger after a predefined number of offline state-action pairs has been used, denoted as $N$. We set $N$ to be 100, 200, or 400 for easier tasks (e.g., Hopper, Ant) and 200, 400, or 800 for harder tasks (e.g., HalfCheetah, Walker2D).
@@
Though the nonrealizable setting is beyond the scope of this work, we expect that some variant of our algorithm can still give reasonable performance, provided that the policy class is expressive enough (so that the approximation error is nonzero but small). For example, [li2023agnostic] observed that with nonrealizable stochastic experts, DAgger variants outperform BC, and exhibit learning curves similar to ours.
@@ G.1 Additional Experiment Plots
We present extended experiment results with larger cost budgets. As shown in Figure 4, we allocate a total annotation budget of 2000 for Hopper and Ant, and 4000 for HalfCheetah and Walker. This complements Figure 3 in the main paper by showing the full training curves without zooming into the stage with small cost budget. The trends are consistent with our earlier observations: Warm-Stagger achieves similar or better sample efficiency compared to Stagger when $C=1$, and clearly outperforms both baselines under the cost-aware setting where $C=2$.
@@ G.2 Experiment with MSE Loss
We additionally evaluate our algorithms using mean squared error (MSE) as the loss function for optimization. All training settings remain identical to the main experiments with log loss, except that we use a learning rate of $2.5\times 10^{-4}$. As shown in Figure 5, we observe qualitatively similar results to those under log loss shown in 3, consistent with prior observations in [foster2024behavior], with the added benefit of more stable training dynamics.
@@
G.3 Additional Experiments with Algorithm 3
For completeness, we evaluate Tragger and its warm-start variant (Warm-Tragger), as shown in Algorithm 4, on continuous control tasks and the same MDP setup as in Figure 2. The key distinction between Warm-Tragger and Warm-Stagger lies in the annotation mode: the former employs trajectory-wise oracle feedback instead of state-wise annotation, leading to notably different behaviors, as shown in Figure 6.
In particular, for Ant and HalfCheetah, the sample efficiency ($C=1$) of Tragger and Warm-Tragger is significantly worse than that of Stagger due to the cold-start problem: early DAgger rollouts have poor state coverage but must still proceed until the end of each trajectory. In contrast, Stagger samples only a single state per trajectory, thereby better leveraging interactive feedback.
For Hopper and Walker, however, Tragger and Warm-Tragger achieve performance closer to Stagger. These environments feature hard resets when the agent fails (unlikely in Ant and never in HalfCheetah), which truncate poor trajectories and consequently improve sample efficiency.
Overall, these observations suggest a natural middle ground between full-trajectory and single-state annotation—namely, batch queries (e.g., sampling 50 states per trajectory), as explored by [li2023agnostic] with comparable results.
@@ G.3 Additional Experiments with Algorithm 3
1:Input: MDP $\mathcal{M}$, trajectory-wise expert annotation oracle $\mathcal{O}^{\mathrm{Traj}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s)=\pi^{\mathrm{E}}(s)\ ,\forall s\in D_{\text{off}}\}$.

### v3

@@ E.4 Proof of Theorem 31
By Lemma 33, for each $\pi^{n}$, which is a first-step mixing of Markovian policies:
$\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\|s_{1:H}),\pi^{\mathrm{E}}(\cdot\|s_{1:H}))\right]\geq\frac{1}{2}\rho(\pi^{n}\parallel\pi^{\mathrm{E}}).$
@@
$\frac{\operatorname{OnEst}^{\mathrm{Traj}}_{N}}{N}=\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\|s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\|s^{n}_{1:H}))\right]\geq\frac{1}{2N}\sum_{n=1}^{N}\rho(\pi^{n}\parallel\pi^{\mathrm{E}})=\frac{1}{2}\rho(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we apply the fact that $\hat{\pi}$ is a first-step uniform mixing of $\{\pi^{n}\}_{n=1}^{N}$. Finally, we conclude the proof by applying Lemma 40. ∎
Proof of Theorem 31.
The proof closely follows Proposition C.2 in [17],which was tailored to another variant of DAgger. Different from that analysis, here we leverage the distribution of the state sequence $s_{1:H}$ instead of the per-step state distribution.
@@
$\ell^{n}(\pi)=\log\left(\frac{1}{\pi(a^{n,*}_{1:H}\parallel s^{n}_{1:H})}\right)$
It can be observed that $\ell^{n}(\pi_{u})$’s are 1-exp-concave in $u\in\Delta(\mathcal{B})$. Therefore, setting $\mathbb{A}$ as the exponential weight algorithm (Proposition 41) ensures that the following bound holds almost surely:
@@
Now, Lemma 42 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}(\cdot\|\cdot)$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s_{1}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\|s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\|s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
The second part of the theorem follows by applying Lemma 37. ∎
@@ Appendix F Auxiliary Results
$\frac{1}{2}\parallel p-q\parallel_{1}\leq D_{H}^{2}(p,q)\leq\parallel p-q\parallel_{1}$
Proof.
Without loss of generality, assume that $q$ is a delta mass on $z_{0}$. Therefore, $\|p-q\|_{1}=2(1-p(z_{0}))$, and
$D_{H}^{2}(p,q)=(1-\sqrt{p(z_{0})})^{2}+(1-p_{z_{0}})=2(1-\sqrt{p(z_{0})}).$
Thus,
$\frac{D_{H}^{2}(p,q)}{\|p-q\|_{1}}=\frac{1}{1+\sqrt{p(z_{0})}}\in[\frac{1}{2},1].\qed$
Lemma 39 (Performance Difference Lemma [23, 49]).
@@
Lemma 40 (Lemma D.2. of  [17]).
@@
Proposition 41 (Proposition 3.1 of [9]).
Suppose $\{\ell^{n}(\cdot)\}_{n=1}^{N}$ is a sequence of $\eta$-exp-concave functions from $\Delta(\mathcal{X})$ to $\mathbb{R}$. For all $x\in\mathcal{X}$, define the weights $w^{n-1}(x)$ and probabilities $u^{n}(x)$ as follows:
$w^{n-1}(x)=e^{-\eta\sum_{i=1}^{n-1}\ell_{i}(e_{x})},\quad u^{n}(x)=\frac{w^{n-1}(x)}{\sum_{x^{\prime}\in\mathcal{X}}w^{n-1}(x^{\prime})},$
where $e_{x}$ is the $x$-th standard basis vector in $\mathbb{R}^{|\mathcal{X}|}$. Then, choosing $u^{n}=(u^{n}(x))_{x\in\mathcal{X}}$ (exponential weights used with learning rate $\eta$) satisfies:
@@
Lemma 42 (Corollary of Lemma A.14 in [18]).
Under the realizbility assumption, where there exists $g_{\star}\in\mathcal{G}$ such that for all $n\in[N]$,
$y^{n}\mid x^{n},\mathcal{H}^{n-1}\sim g_{\star}(\cdot\mid x^{n}),$
where $\mathcal{H}^{n-1}$ denotes all histories at the beginning of round $n$. Then, for any estimation algorithm that outputs $\left\{\hat{g}^{n}\right\}_{n=1}^{N}$ online and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sum_{n=1}^{N}\mathbb{E}_{n-1}\left[D_{\mathcal{H}}^{2}\left(\hat{g}^{n}(\cdot\mid x^{n}),g_{\star}(\cdot\mid x^{n})\right)\right]\leq\sum_{n=1}^{N}\left(\log\frac{1}{\hat{g}^{n}(y^{n}\mid x^{n})}-\log\frac{1}{g_{\star}(y^{n}\mid x^{n})}\right)+2\log(\delta^{-1}).$
where $\mathbb{E}_{n}[\cdot]:=\mathbb{E}[\cdot\mid\mathcal{H}^{n}]$.
@@ Appendix G Experiment Details
We compare Warm-Stagger with Behavior Cloning (BC) and Stagger on continuous-control tasks from OpenAI Gym MuJoCo [72, 8] with episode length $H=1000$.
@@
All experiments were conducted on a Linux workstation equipped with an Intel Core i9 CPU (3.3GHz) and four NVIDIA GeForce RTX 2080 Ti GPUs. Our implementation builds on the publicly available DRIL repository [7] (https://github.com/xkianteb/dril), with modifications to support interactive learning. The continuous control environments used in our experiments are: “HalfCheetahBulletEnv-v0”, “AntBulletEnv-v0”, “Walker2DBulletEnv-v0”, and “HopperBulletEnv-v0”. We include link to our implementation here: https://github.com/liyichen1998/Interactive-and-Hybrid-Imitation-Learning-Provably-Beating-Behavior-Cloning.
@@
We use four MuJoCo environments: Ant, Hopper, HalfCheetah, and Walker2D. The expert policy is a deterministic MLP pretrained via TRPO [52, 53], with two hidden layers of size 64.
Model Architecture used by Learner.
The learner uses the same MLP architecture as the expert. Following [17], we use a diagonal Gaussian policy:
@@
To evaluate the performance of BC against the number of states annotated, we reveal expert state-action pairs sequentially along expert trajectories until the annotation budget is reached. For Stagger, at each iteration, it rolls out the latest policy, samples a state uniformly from the trajectory, queries it for the expert action, and updates immediately.
For Warm-Stagger, we begin with BC and switch to Stagger after a predefined number of offline expert state-action pairs has been used, denoted as $N$. We set $N$ to be 100, 200, or 400 for easier tasks (e.g., Hopper, Ant) and 200, 400, or 800 for harder tasks (e.g., HalfCheetah, Walker2D).
@@
Though the nonrealizable setting is beyond the scope of this work, we expect that some variant of our algorithm can still give reasonable performance, provided that the policy class is expressive enough (so that the approximation error is nonzero but small). For example, [34] observed that with nonrealizable stochastic experts, DAgger variants outperform BC, and exhibit learning curves similar to ours.
@@ G.1 Additional Experiment Plots
We present extended experiment results with larger cost budgets. As shown in Figure 4, we allocate a total annotation cost budget of 2000 for Hopper and Ant, and 4000 for HalfCheetah and Walker. This complements Figure 3 in the main paper by showing the full training curves without zooming into the stage with small cost budget. The trends are consistent with our earlier observations: Warm-Stagger achieves similar or better sample efficiency compared to Stagger when $C=1$, and clearly outperforms both baselines under the cost-aware setting where $C=2$.
@@ G.2 Experiment with MSE Loss
We additionally evaluate our algorithms using mean squared error (MSE) as the loss function for optimization. All training settings remain identical to the main experiments with log loss, except that we use a learning rate of $2.5\times 10^{-4}$. As shown in Figure 5, we observe qualitatively similar results to those under log loss shown in Figure 3, consistent with prior observations in [17], with perhaps more stable learning curves.
@@
G.3 Additional Experiments with Tragger
For completeness, we evaluate Tragger and its warm-start variant (Warm-Tragger, Algorithm 4), on continuous control tasks and the toy MDP as in Figure 2. The key distinction between Warm-Tragger and Warm-Stagger lies in the annotation mode: the former employs trajectory-wise oracle feedback instead of state-wise annotation, leading to notably different behaviors, as shown in Figure 6.
In particular, for Ant and HalfCheetah, the state-wise cost efficiency ($C=1$) of Tragger and Warm-Tragger is significantly worse than that of Stagger due to the cold-start problem: early DAgger rollouts have poor state coverage but must still proceed until the end of each trajectory. In contrast, Stagger samples only one state per trajectory, thereby better leveraging interaction to get timely feedback. For Hopper and Walker, however, Tragger and Warm-Tragger achieve performance closer to Stagger. These environments feature hard resets when the agent’s state becomes unhealthy (unlikely in Ant and never in HalfCheetah), which terminates (’truncate’ has a specific meaning of ending at a prespecified step in openai gym) poor trajectories and consequently improve sample efficiency.
Overall, these observations suggest a natural middle ground between full-trajectory and single-state annotation—namely, batch queries (e.g., sampling 50 states per trajectory), as explored by [34] with comparable results.
@@ G.3 Additional Experiments with Tragger
1:Input: MDP $\mathcal{M}$, trajectory-wise expert annotation oracle $\mathcal{O}^{\mathrm{Traj}}$, Stationary policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\mathrm{off}}$ of size $N_{\mathrm{off}}$, interaction budget (in terms of number of states) $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s_{h})=a_{h},\ \forall h\in[H],\forall(s,a)_{1:H}\in D_{\text{off}}\}$